OpenStudy (anonymous):

Judah makes pottery to sell at his school’s craft bazaar. Every year, he determined that he makes twenty-five dollars less than he made the previous year. Write a function rule to define the total amount y Judah made based on the amount x he made the previous year. Give an example of how the function rule works. Which is the dependent variable? The independent variable?

OpenStudy (anonymous):

medal!!

OpenStudy (anonymous):

For every year he makes 25 less than the last year. If x was the amount he made the previous year It would be y=x-25. to give an example, say he made 50 last year he would make y=50-25=25 this year The dependent variable is the one that relies on the other since you input x it is in the independent and the y is the dependent

OpenStudy (anonymous):

ooohhhh ok!!!! thanks!!!

OpenStudy (anonymous):

no problem